# The Influence of Human-inspired Agentic Sophistication in LLM-driven Strategic Reasoners

### v1

@@ 2 Human Reasoning in 2-Player Guessing Games
A solution in guessing games is typically modelled using k-level reasoning theory [6], which assumes that each level-$k$ player believes others reason at most at level $k-1$ and best responds accordingly. This recursive process is formalised as: $a_{k}=p^{k}\cdot a_{0}$, where $a_{0}$ denotes the choice of 0-level players and $a_{k}$ the response of level-$k$ players. While in $n$-player repeated games players often iteratively reject weakly dominant strategies, gradually driving the average guess toward zero, the dynamics change considerably in two-player settings. Specifically, when $p<1$, choosing 00 becomes a strictly dominant strategy, as the lower guess will always win. Despite this theoretical simplicity, empirical studies show that players frequently select dominated strategies [28].
@@
Theoretical expectations suggest that the expert group is more likely to identify the dominant strategy than students. This is reflected in our dataset, as $11.9\%$ of students and $37.7\%$ of experts chose 00. Furthermore, Figure 1 supports the expected intergroup performance difference: the expert cohort displays a substantially lower mean guess and greater density over lower values. Experts also exhibit higher skewness ($1.50$ vs. $0.55$). A Levene’s test found no significant difference in variances between the groups ($p=0.4351$), justifying the use of a standard independent samples $t$-test, which confirms a significant difference at $t=4.04$, $p=0.0001$.
@@ 3.1 Multi-agent Simulation Framework
A centralised multi-agent framework for agentic reasoning was introduced [38], combining an Observe–Orient–Decide–Act (OODA) control loop [39] with multi-level hypergames [43] to support recursive reasoning among agents. Here we define a streamlined variant that omits explicit evaluation of recursive reasoning and focuses on core agent interactions within our domain. We let $X$ denote the space of natural language game descriptions and $C$ the set of agent contexts. An umpire $\upsilon$, conceptualised as a pseudo-player [33], initialises each agent with a context $c_{i}\in C$ and instruction model $m_{i}\in M$. The umpire further enables gameplay by pairing agents and managing their participation in the guessing game, denoted as $G^{*}$.
Players’ internal processes are defined as the interpretation function $I:X\times C\times M\rightarrow\mathbb{Z}$ – wrapping the agent cycle’s orient and decision stages –, which we refer to as a cognitive process, producing a guess from a natural language game description, player context and instruction. Our experiments implement agents with different implementations of $I$.
The work in [38] entails a perspective-based formal definition of guessing games that captures individuals’ levels of recursive thinking. Our 2-player guessing games in this context are defined as $G=(N,p,A,U)$, where $N=\{i,j\}$ are our two players, and $p=\frac{2}{3}$ is the target multiplier. The action space is $A=A_{i}=A_{-i}\subseteq\mathbb{Z}$, where $i$ and $-i$ are the player and their opponent, $A$ is a finite interval of integers shared by both players, such as $A=\{0,1,\dotsc,100\}$ and $U:A\rightarrow\mathbb{R}$ is the utility function, specified as Equation 1 below.
$U_{i}(a_{i},a_{-i})=\begin{cases}1&\text{if }|a_{i}-p\mu|<|a_{-i}-p\mu|\\ 0.5&\text{if }|a_{i}-p\mu|=|a_{-i}-p\mu|\\ 0&\text{otherwise}\end{cases}$ (1)
where $\mu=\frac{a_{i}+a_{-i}}{2}$.
In the following sections, we introduce a set of agents within the framework defined above. First, we consider a game-theoretic model as a simple benchmark agent, then present two LLM-driven artificial reasoners that implement $I$ with different levels of conceptual sophistication.
@@ 3.2 Experience Weighted Attraction
Camerer et al. introduced the self-tuning experience weighted attraction model (EWA) as a single free-parameter game-theoretic learning algorithm designed to produce results that mimic human behaviour [16]. We implement the self-tuning EWA as a benchmark, given that it is a well-established state-of-the-art model from the economics literature. The heart of the model is the attraction mechanism composed as follows.
$\text{Memory}^{j}_{i}(t)=\phi\cdot N(t-1)\cdot\mathcal{A}^{j}_{i}(t-1)$ (2)
The memory component defines how past experience is carried forward, composed of the experience weight $N$ [5] and the change detector function $\phi$, defined in equations 3 and 4 below.
$N(t)=(1-\kappa)\cdot\phi\cdot N(t-1)+1$ (3)
$\phi(t)=1-\frac{1}{2}S_{i}(t)$ (4)
The surprise index function is composed of the immediate change vector $h$ and cumulative change vector $r$:
$S_{i}(t)=\sum^{m_{-i}}_{k=1}{(h^{k}_{i}(t)-r^{k}_{i}(t))^{2}}$ (5)
$h^{k}_{i}(t)=\frac{\sum^{t}_{\tau=1}{\mathcal{I}(s^{k}_{-i},s_{-i}(\tau))}}{t}$ (6)
$r^{k}_{i}(t)=\mathcal{I}(s^{k}_{-i},s_{-i}(t))$ (7)
where $\mathcal{I}$ is a binary index function.
The reinforcement component describes how new experiences are recorded:
$\text{Reinforcement}^{j}_{i}(t)=[\delta+(1-\delta)\cdot\mathcal{I}(s^{j}_{i},s_{i}(t))]\cdot\pi_{i}(s^{j}_{i},s_{-i}(t))$ (8)
The foregone payoffs are denoted by $\delta$:
@@
where $\pi_{i}(s^{j}_{i},s_{-i}(t))$ is $i^{\prime}s$ payoff for selecting strategy $j$ against $s_{-i}$ at time $t$. Players’ attractions are updated according to equation 10 below
$\mathcal{A}^{j}_{i}(t)=\frac{\text{Memory}^{j}_{i}(t)+\text{Reinforcement}^{j}_{i}(t)}{N(t-1)\cdot\phi\cdot(1-\kappa)}$ (10)
where $\kappa=0$, and players choose a strategy according to the distribution $P^{j}_{i}(t+1)=\frac{e^{\lambda\cdot\mathcal{A}^{j}_{i}(t)}}{\sum^{m_{i}}_{k=1}{e^{\lambda\cdot\mathcal{A}^{k}_{i}(t)}}}$, where $\lambda$ is the single parameter that requires manual tuning and controls the sensitivity of the model. The initial attraction $\mathcal{A}^{j}_{i}(0)$ is implemented according to the cognitive hierarchy model and the Poisson distribution function $P(k)=\frac{e^{-\tau}\tau^{k}}{k!}$ [6], and we set $N(0)=1$. Finally, we adopt $\lambda=2.39$ and $\tau=1.5$, which are the reported optimal parameters, tested on n-player guessing games [5].
Given that the self-tuning EWA model does not process natural language, we integrate EWA with our framework by implementing $I_{\text{EWA}}:G\rightarrow\mathbb{Z}$, where $I_{\text{EWA}}(G)\sim P(t+1)$ denotes sampling from the induced probability distribution over the strategies. In addition, in the case of EWA players, the responsibilities of $\upsilon$ are extended by the translation function $T:X\rightarrow G$, and the subject of game requests is $G^{*}$ in place of the natural language description $X$.
@@ 4.2.1 Population-wise Performance
First, we consider the case where the players’ individual details – i.e. the personal profile defining whether the player is a student or an expert – are not relevant. Data with subgroup representation is pooled together to represent the overall population.
In order to estimate and discuss the reasoning sophistication of our different implementations, we convert the guess values to k-levels as follows. Assuming irrational $k=0$ actors act randomly and produce a mean of 50 [6], we solve $(\frac{2}{3})^{k}\cdot 50=a_{i}^{*}$ to obtain $i$’s reasoning level. We set one dimension as agentic sophistication, where the order of sophistication is set as $\text{EWA}<S<R$. The second dimension describes the sophistication of the implemented $C$ and $M$, such that
$c_{\varnothing},m_{0}<c_{\varnothing},m_{1}<c_{\text{sim}},m_{0}<c_{\text{sim}},m_{1}<c_{\text{bio}},m_{0}<c_{\text{bio}},m_{1}$
defines the order of sophistication. Our initial expectation is that there is a linear relation between the two axes of sophistication and the human-likedness of the results. In Figure 3, we present LLM-wise heatmaps of the absolute mean error (AME) $\left|\bar{k}_{\text{true}}-\bar{k}_{\text{pred}}\right|$ of the true human and the agent predicted k values. We can observe that the least complex implementation (EWA) showed the most distant performance, and Haiku produced the best results with the most complex implementation, which matches humans’ reasoning level with a difference of just $0.02$. However, while the opposite ends of the complexity range matched our expectations, the intermediate results are not always consistent with the progressive inclusion of more sophisticated implementations. Sonnet shows the opposite reaction to more advanced configurations and, overall, is more distant from human reasoners than the simpler LLM.
We transform our categorical features into ordinal variables as follows: agent $\in\{\text{EWA}=0,S=1,R=2\}$, $c\in\{c_{\varnothing}=0,c_{sim}=1,c_{bio}=2\}$, $m\in\{m_{0}=0,m_{1}=1\}$, and model $\in\{\text{Haiku}=0,\text{Sonnet}=1\}$. Then, the overall sophistication score is computed as their sum. Table 1 reveals the statistical effects of individual sophistication factors on our LLM-based agents, measured by linear regression models on the k-level AME. While the agent type does not significantly influence the results, the sophistication of LLM, $C$ and $M$ have significant effects. However, the effect direction is not aligned with our expectations, and a Spearman correlation test confirms a weak and insignificant negative correlation between the sophistication score and deviation from humans’ k-level.
@@
While the corresponding k-levels are convenient and meaningful measures of the overall performance, they are not informative enough for individual-based simulations where the goal is to match the distribution of human guesses. We calculate the kernel density estimates and use the Wasserstein distance to measure the similarity between the distributions of human and agent guesses. The distributions are presented in Figure 4. Table 2 summarises the Wasserstein distances between each agent configuration per LLM and human guesses and Table 1 contains the statistical effects of sophistication factors on the distance. The distribution-related results paint a slightly different picture than the population mean-derived metrics. While, overall, Haiku still outperforms Sonnet in matching human results, and the best match is achieved by $R$, the most human-like result is produced by the least sophisticated $c_{\varnothing},m_{o}$ configuration. Similarly to before, while Haiku benefits from the more sophisticated agent implementation, Sonnet performs better with the simpler alternative. Another key difference between the two models is that the inclusion of MoA had a positive effect on Haiku in most configurations with both $S$ and $R$, but it had little to no effect on Sonnet’s reasoning.
Figure 4: KDEs of $S$ (top) and $R$ (bottom) agents with $C$,$M$ variations. Red and blue correspond to overall results, light/dark shades correspond to cohorts.
@@
Config EWA Simple Reasoner
@@
$c_{\varnothing},m_{0}$ 22.34 22.05 18.59 14.28* 20.38
$c_{\varnothing},m_{1}$ N.A. 16.69 19.79 17.51 20.88
$c_{\text{sim}},m_{0}$ N.A. 19.92 15.41 17.46 19.32
$c_{\text{sim}},m_{1}$ N.A. 15.05 14.38 17.42 18.88
$c_{\text{bio}},m_{0}$ N.A. 22.99 16.57 18.37 20.33
$c_{\text{bio}},m_{1}$ N.A. 19.08 16.24 16.14 20.78
The human data indicates an overall rate of $23.15\%$ of players managed to pick the equilibrium strategy: 00. While EWA produced the highest mean k-level, indicating a bias towards optimisation instead of human-level performance, it still fails to derive the dominant strategy. In the case of the LLM-based agents, our general observation so far was that game-theoretic problem solving is more dominant in Sonnet’s results than in Haiku’s. While by the previous metrics, this had a negative effect on Sonnet’s ability to replicate human results, the current results show it helps with reaching zero guesses – shown in Table 3 – and zero choices appear in most settings. Haiku is unable to replicate zero guesses in most cases, with $S_{c_{\text{sim},m_{1}}}$ as the sole exception at $14\%$. While MoA improves the rate of 00 guesses for $S$ with Sonnet and both models’ best results are associated with MoA, the overall sophistication score does not significantly correlate with the agents’ performance of matching humans’ 00 guessing rate.
Table 3: 00 guesses per 100 agents (Sonnet only). Human benchmark at $23$.
@@ 4.2.2 Subpopulation-wise Performance
Simulations that require an individual-based approach where specific subpopulations can be represented explicitly must implement role specification. The following analysis concerns the configurations supporting profile assignment and investigates how well artificial agents can replicate the intergroup differences, role-specific reasoning levels, and guess distributions. Figure 1 depicting human results per cohort shows a higher density of guesses over lower values and a significantly lower mean of guesses for experts. In this section, we use those results as a benchmark and evaluate artificial agents’ ability to replicate the significant difference in reasoning between human students and experts.
First, we calculate the k-level means for students and experts separately and define $\Delta=\left|\bar{k}_{\text{true}}^{\text{student}}-\bar{k}_{\text{pred}}^{\text{student}}\right|-\left|\bar{k}_{\text{true}}^{\text{expert}}-\bar{k}_{\text{pred}}^{\text{expert}}\right|$ to measure if the agent replicates student and expert behaviour equally well/poorly, where values close to 00 indicate an approximation of the difference between the human groups. Figure 5 entails the $\Delta$ heatmaps for Haiku and Sonnet. The dominantly negative values in the case of $S$ agents across both models indicate that these agents are consistently closer to expert behaviour, where likely the overestimation of student reasoning levels is due to an overfit to game-theoretic optimizations we observed in previous instances. On the other hand, $R$ agents show a more balanced behaviour and manage to approximate better the human differences with all $C\times M$ errors within +/0.4+/-0.4+ / - 0.4 k levels. In addition, the transition from $m_{0}$ to $m_{1}$ consistently causes agents to optimise more and overestimate reasoning levels. The statistical significance test indicates lesser divergence in the case of $R$ agents using Haiku and the minor and insignificant differences in the case of $R_{c_{\text{sim}},m_{0}}$ and $R_{c_{\text{bio}},m_{0}}$ indicate closer matching of the human intergroup divergence.
Figure 5: Delta between agents’ and humans’ cohort-wise mean k-level difference with significance levels: * p < 0.05, ** p < 0.01, *** p < 0.001.
Figure 4 is extended with the profile-specific distribution curve per model for the configurations where the distinction between cohorts is available, and Table 4 collects the Wasserstein distances for each agent variation.

### v2

@@ 2 Human Reasoning in 2-Player Guessing Games
A solution in guessing games is typically modelled using k-level reasoning theory [6], which assumes that each level-$k$ player believes others reason at most at level k1k-1 and best responds accordingly. This recursive process is formalised as: $a_{k}=p^{k}\cdot a_{0}$, where $a_{0}$ denotes the choice of 0-level players and $a_{k}$ the response of level-$k$ players. While in $n$-player repeated games players often iteratively reject weakly dominant strategies, gradually driving the average guess toward zero, the dynamics change considerably in two-player settings. Specifically, when $p<1$, choosing 0 becomes a strictly dominant strategy, as the lower guess will always win. Despite this theoretical simplicity, empirical studies show that players frequently select dominated strategies [28].
@@
Theoretical expectations suggest that the expert group is more likely to identify the dominant strategy than students. This is reflected in our dataset, as $11.9\%$ of students and $37.7\%$ of experts chose 0. Furthermore, Figure 1 supports the expected intergroup performance difference: the expert cohort displays a substantially lower mean guess and greater density over lower values. Experts also exhibit higher skewness ($1.50$ vs. $0.55$). A Levene’s test found no significant difference in variances between the groups ($p=0.4351$), justifying the use of a standard independent samples $t$-test. The test confirms that the difference in means is statistically significant ($t=4.04$, $p=0.0001$).
@@ 3.1 Multi-agent Simulation Framework
We build on our earlier work [38], which introduced a centralised multi-agent framework combining the Observe–Orient–Decide–Act (OODA) control loop [39] with multi-level hypergames [43] to support recursive reasoning. For the present study, we define a streamlined variant that omits explicit recursive evaluation and instead emphasises the core interaction cycle between agents.
Framework Overview
The framework centres on an umpire $\upsilon$, conceptualised as a pseudo-player [33]. The umpire initialises each player with a context $c_{i}\in C$ and an instruction model $m_{i}\in M$, where $C$ is the set of agent contexts and $M$ the set of instruction templates. The umpire also manages gameplay, pairing agents and orchestrating their participation in the guessing game.
Agent Processes
Given a natural language description $x\in X$, where $X$ is the space of game descriptions, each player $i$ applies an interpretation function:
$I:X\times C\times M\rightarrow\mathbb{Z},$
which encapsulates the orient and decide phases of the agent cycle. We refer to $I$ as the player’s cognitive process, mapping $(x,c_{i},m_{i})$ to a numerical guess. Different implementations of $I$ capture distinct forms of reasoning, and their comparison constitutes the focus of our study on agentic sophistication.
Game Formalisation
Following [38], we adopt a perspective-based formalisation of two-player guessing games. A game is defined as
$G=(N,p,A,U),$
where $N=\{i,j\}$ are the two players, $p=\tfrac{2}{3}$ is the target multiplier, and $A=A_{i}=A_{-i}\subseteq\mathbb{Z}$ is a shared action space, typically $A=\{0,1,\dots,100\}$. The utility function $U:A\times A\to\mathbb{R}$ is
$U_{i}(a_{i},a_{-i})=\begin{cases}1&\text{if }|a_{i}-p\mu|<|a_{-i}-p\mu|,\\ 0.5&\text{if }|a_{i}-p\mu|=|a_{-i}-p\mu|,\\ 0&\text{otherwise},\end{cases}$ (1)
where $\mu=\tfrac{a_{i}+a_{-i}}{2}$ is the mean guess.
Roadmap
In the following sections, we instantiate this framework with several agent types. We begin with a game-theoretic benchmark (EWA), then introduce two LLM-based artificial reasoners, each implementing $I$ with different levels of conceptual sophistication.
@@ 3.2 Experience Weighted Attraction
We choose as a benchmark the self-tuning experience weighted attraction model (EWA) as a single-parameter game-theoretic learning algorithm designed to produce results that mimic human behaviour [16]. The model relies on three specific modeling features that assign attraction values to the agent’s strategies, which are (i) reinforced by earned payoffs, (ii) simulated by considering foregone payoffs, and (iii) decayed by a memory factor. The model is driven by the update function $\mathcal{A}_{j}^{i}(t)$ assigning attention values at each step $t$ to a player $i$’ strategy $j$ – denoted by $s_{i}^{j}$ –, defined as follows:
$A^{j}_{i}(t)=\frac{\text{Memory}^{j}_{i}(t)+\text{Reinforce}^{j}_{i}(t)}{N(t-1)\cdot\phi(t)\cdot(1-\kappa)}.$ (2)
The Memory component integrates a cumulative experience weight $N$ [5] and defines how past experience is carried forward:
$\text{Memory}^{j}_{i}(t)=\phi(t)\cdot N(t-1)\cdot\mathcal{A}^{j}_{i}(t-1),$ (3)
where $\phi$ denotes a change detector function
$\phi(t)=1-\frac{1}{2}S_{i}(t),$ (4)
specifying how much the agent’s current observation $h$ differs from the accumulated past experience $r$, encapsulated in the Surprise index $S_{i}$:
$S_{i}(t)=\sum^{m_{-i}}_{k=1}{(h^{k}_{i}(t)-r^{k}_{i}(t))^{2}}.$ (5)
Given player payoff $\pi_{i}$, EWA’s choice reinforcement is modeled as:
$\text{Reinforce}^{j}_{i}(t)=[\delta+(1-\delta)\cdot\mathcal{I}(s^{j}_{i},s_{i}(t))]\cdot\pi_{i}(s^{j}_{i},s_{-i}(t)),$ (6)
where $\delta$ denotes the simulation of “would have earned” payoffs:
@@
Finally, A softmax exponential form prediction function is used over attractions to sample next period play, leaving $\lambda$ – controlling sensitivity – as a single free parameter to be tuned:
$P^{j}_{i}(t+1)=\frac{e^{\lambda\cdot\mathcal{A}^{j}_{i}(t)}}{\sum^{m_{i}}_{k=1}{e^{\lambda\cdot\mathcal{A}^{k}_{i}(t)}}}.$ (8)
[…]
Since EWA operates directly on formal game representations rather than natural language, we integrate it into our framework by defining a dedicated interpretation function
$I_{\text{EWA}}:G\rightarrow\mathbb{Z},\qquad I_{\text{EWA}}(G)\sim P_{i}(t{+}1),$
which samples a strategy according to the softmax distribution in Eq. (8). In this case, the umpire $\upsilon$ is responsible for translating natural language descriptions $x\in X$ into their corresponding formal game representations via a mapping $T:X\rightarrow G$. Accordingly, when EWA agents participate in the framework, the subject of their game requests is the formalised game rather than the natural language description.
@@ 4.2.1 Population-wise Performance
We first analyse performance at the population level, pooling data across subgroups to represent the overall population of participants. This evaluation captures how well agents approximate the aggregate patterns of human reasoning.
Reasoning Sophistication
We convert numerical guesses to $k$-levels, allowing us to measure performance in terms of reasoning depth. Following [6], $k=0$ actors are assumed to act randomly with a mean guess of 50, and reasoning depth is derived by solving
$\left(\tfrac{2}{3}\right)^{k}\cdot 50=a_{i}^{*}.$
We treat architecture as the first dimension of sophistication in increasing order: $\text{EWA}<S<R$. Secondly, the sophistication of context/instruction configurations $(C,M)$ is given by the ordering:
$c_{\varnothing},m_{0}<c_{\varnothing},m_{1}<c_{\text{sim}},m_{0}<c_{\text{sim}},m_{1}<c_{\text{bio}},m_{0}<c_{\text{bio}},m_{1}.$
Our initial expectation was that increasing sophistication along both axes would monotonically improve alignment with human behaviour.
Figure 3 shows LLM-wise heatmaps of the absolute mean error (AME) $\lvert\bar{k}_{\text{true}}-\bar{k}_{\text{pred}}\rvert$. At the extremes, results align with intuition: EWA is the most distant, while Haiku, with the most elaborate configuration, matches humans almost exactly (AME =0.02=0.02). However, intermediate cases deviate from this trend: Sonnet, in particular, often performs worse with more sophisticated configurations, and overall diverges more from human reasoners than Haiku.
Factor Analysis
We transform our categorical features into ordinal variables as follows: agent $\in\{\text{EWA}=0,S=1,R=2\}$, $c\in\{c_{\varnothing}=0,c_{sim}=1,c_{bio}=2\}$, $m\in\{m_{0}=0,m_{1}=1\}$, and model $\in\{\text{Haiku}=0,\text{Sonnet}=1\}$. Then, the overall sophistication score is computed as the sum of their scores. Table 1 illustrates the statistical impact of individual factors on our LLM-based agents, as measured by linear regression models applied to the k-level AME. While the type of agent does not significantly affect the results, $C$ and $M$ show a significant impact. Interestingly, the direction of this effect does not align with our initial expectations. Additionally, a Spearman correlation test reveals a weak and statistically insignificant negative correlation between the sophistication score and the deviation from human k-level performance.
@@
Distributional Similarity
Since $k$-levels summarise only mean behaviour, we also assess complete guess distributions using kernel density estimates (Figure 4) and Wasserstein distance. Tables 2 and 1 report the distributional distances and the statistical effects of individual factors. Haiku again generally outperforms Sonnet, and Reasoner ($R$) agents achieve the closest match. Nevertheless, the most human-like outcome is produced by the least sophisticated configuration ($c_{\varnothing},m_{0}$), highlighting that greater sophistication does not guarantee better alignment. Another difference between the LLMs is that MoA consistently benefits Haiku but has a negligible effect on Sonnet.
Figure 4: KDEs of $S$ (top) and $R$ (bottom) agents with $C$,$M$ variations. Red and blue correspond to overall results, and lighter/darker shades correspond to specific cohorts.
@@
Config
EWA Simple Reasoner
@@
$c_{\varnothing},m_{0}$
22.34 22.05 18.59 14.28* 20.38
$c_{\varnothing},m_{1}$
N.A. 16.69 19.79 17.51 20.88
$c_{\text{sim}},m_{0}$
N.A. 19.92 15.41 17.46 19.32
$c_{\text{sim}},m_{1}$
N.A. 15.05 14.38 17.42 18.88
$c_{\text{bio}},m_{0}$
N.A. 22.99 16.57 18.37 20.33
$c_{\text{bio}},m_{1}$
N.A. 19.08 16.24 16.14 20.78
Equilibrium Play
Finally, we evaluate agents on their ability to produce the equilibrium strategy: picking 0. Human players achieve this in $23.15\%$ of cases. EWA, despite producing higher $k$-levels than humans, never selects zero. As Table 3 reveals, Sonnet agents generate zero guesses in most settings, with MoA boosting rates substantially, whereas Haiku rarely produces zeros – the sole exception being $S_{c_{\text{sim}},m_{1}}$ at $14\%$. Overall, zero-guessing behaviour does not significantly correlate with the sophistication score.
Table 3: 0 guesses per 100 agents (Sonnet only). Human benchmark at $23$.
@@ 4.2.2 Subpopulation-wise Performance
Simulations that require an individual-based approach, where specific subpopulations can be represented explicitly, must implement role specification. The following analysis concerns the configurations supporting profile assignment and investigates how well artificial agents can replicate the intergroup differences, role-specific reasoning levels, and guess distributions. As shown in Figure 1, human experts exhibit a substantially lower mean guess and greater density over lower values than students. These results provide the benchmark for our subpopulation-level evaluation.
Reasoning Sophistication
First, we calculate the k-level means for students and experts separately and define $\Delta=\left|\bar{k}_{\text{true}}^{\text{student}}-\bar{k}_{\text{pred}}^{\text{student}}\right|-\left|\bar{k}_{\text{true}}^{\text{expert}}-\bar{k}_{\text{pred}}^{\text{expert}}\right|$ to measure whether an agent approximates both groups equally well, with values close to zero indicating better replication of intergroup differences. Figure 5 entails the $\Delta$ heatmaps for Haiku and Sonnet. The dominantly negative values in the case of $S$ agents across both models indicate that these agents are consistently closer to expert behaviour, likely due to an overestimation of student reasoning levels, which is likely an overfit to game-theoretic optimizations we observed in previous instances. In contrast, $R$ agents show more balanced behaviour, with errors within $\pm 0.4$ $k$-levels across all $C\times M$ settings. Transitioning from $m_{0}$ to $m_{1}$ consistently increases optimisation, raising predicted reasoning levels. Statistical tests confirm that divergence is significantly lower for $R$ agents using Haiku, with particularly close Delta matches in $R_{c_{\text{sim}},m_{0}}$ and $R_{c_{\text{bio}},m_{0}}$.
Figure 5: Delta error between agents’ and humans’ mean k-level difference for students and experts with significance levels: * p < 0.05, ** p < 0.01, *** p < 0.001.
Distributional Similarity
We extend the analysis to distributional alignment, adding profile-specific KDE curves to Figure 4 and reporting Wasserstein distances in Table 4. For Haiku, the inclusion of MoA correlates with lower distances, whereas Sonnet exhibits inconsistent patterns without clear trends. The relative overoptimization of Sonnet and its poorer instruction following ability leads to a similar resistance to increasing complexity as we have seen in the experiments without profile specification. In three of four cases, Haiku paired with $c_{\text{sim}},m_{1}$ achieves the closest match. In the case where Haiku did worse, it produced an outlier distribution with unusually low variance and a concentration around $k=1$ and $k=2$, suggesting that the model latched onto canonical game-theoretic reasoning patterns. While a manual inspection of the produced reasoning samples supports this hypothesis, a deeper semantic analysis of reasoning traces lies outside the scope of this paper.